What is the amount of power applied by the wind glass to a 20-kg 

anchor if it is lifted 2-m high with 5s?


1
Expert's answer
2021-01-14T10:40:22-0500

"P=\frac{W}{t}=\frac{mgh}{t}=\frac{20\cdot9.81\cdot2}{5}=78.48(W)" . Answer